X Y
A 24 31
B 13 66
2. What is P(X)?
3. What is P(X or Y)?
4. What is P(X or A)?
5. What is P(X|A)?

Answer by ewatrrr(24785)

 
Hi
always recommend completing table with Totals:
X Y Tot
A 24 31 55
B 13 66 79
Tot 37 97 134
.........
P(X) = 37/134
What is P(X or Y) = 1
P(X or A) = 55/134
P(X|A) = 24/134